# BLEND: Behavior-guided Neural Population Dynamics Modeling via Privileged Knowledge Distillation

### v1

@@ 1 Introduction
To this end, in this paper, we propose BLEND, the Behavior-guided neuraL population dynamics modElling framework via privileged kNowledge Distillation. As shown in Fig. 1(c), BLEND constitutes a student and a teacher model, in which the teacher trains on both behavior observations and neural activity recordings, then distills knowledge to guide the student which takes only neural activity as input. This ensures the student model can make predictions using only recorded neural activity during the deployment/inference stage, but also benefits from the guidance from behavior information during the training stage, making it more versatile for settings lacking behavioral data. Our main contributions are summarized as follows:
@@
Figure 2: Illustration of the proposed BLEND framework, exemplified using neural spiking activity data. (a) Behavior-guided neural representation learning via privileged knowledge distillation. The teacher model is trained on a composite of neural activity and behavioral signals, subsequently distilling its knowledge to a student model that utilizes solely neural activity as input. (b) During the inference phase, the distilled student model is employed for neural population activity analysis and transcriptomic identity prediction tasks.
@@ 2 Related work
Neural dynamics modeling (NDM). NDM refers to a category of methods that aim to capture the dynamics of neural activity by using the activity recordings as inputs. A common and effective choice is the latent variable model (LVM), which leverages low-dimensional latent factors to interpret these dynamics. Various LVMs are developed, ranging from simple non-temporal models such as principal components analysis (PCA) (Cunningham & Yu, 2014) and its variants (Kobak et al., 2016), to linear dynamical systems (Macke et al., 2011; Gao et al., 2016), and to complex state space models like LFADS (Pandarinath et al., 2018). Since the advent of Transformer, its ability to capture temporal dependencies and long-range interactions makes it appropriate for neural data, with representative works such as NeuralDataTransformer (NDT) (Ye & Pandarinath, 2021), STNDT (Le & Shlizerman, 2022), and EIT (Liu et al., 2022). As shown in Fig. 1(a), methods in this category purely depend on neural activity recordings for neural population dynamics modeling yet ignore using the paired behavior information as guidance.
Neural dynamics modeling with behavior as a prior. In recent years, an emerging research direction has focused on jointly modeling neural population dynamics and behavioral signals. As illustrated in Fig. 1(b), these methods can be categorized into two types. The first inputs both behavior and neural activity and utilizes behavior signals to guide the learning of neural dynamics: pi-VAE (Zhou & Wei, 2020) considers the behavior variables as constraints for the construction of latent space of LVMs; CEBRA (Schneider et al., 2023b) utilizes behavior signals to construct contrastive learning samples for label-informed neural activity analysis. However, these approaches normally need to develop delicately designed modules or complex training strategies to achieve their goal. The second line of work aims to decompose neural activity into behavior-relevant and behavior-irrelevant dynamics and reconstruct both behavior signals and neural activity signals. To that end, a linear state-space model, PSID (Sani et al., 2021), is developed to decode population dynamics from motor brain regions. Nonlinear state-space models such as TNDM (Hurwitz et al., 2021) and SABLE (Jude et al., 2022) are proposed to further capture the nonlinear dynamics of neural activity and behavior observation. Nonetheless, these models assume a clear distinction between behaviorally relevant and irrelevant dynamics in input neural activity, which might not always be practical, potentially leading to oversimplification.
In contrast, our approach offers a model-agnostic learning paradigm that can be directly applied to existing NVM models without relying on strong assumptions, thereby circumventing the issues present in current NVM models that integrate behavior information as a prior.
@@ 3 Methods
This section introduces the details of our proposed BLEND framework, i.e., the behavior-guided neural population dynamics modeling via privileged knowledge distillation. We start with problem formulation of behavior-guided neural dynamics modeling in Section 3.1, followed by an exposition of the BLEND algorithm in Section 3.2.
@@ 3.1 Problem Formulation
We concretize the learning problem in this study with neural spiking data. For each trial of the input neural activity, let $\mathbf{x}\in\mathcal{X}=\mathbb{N}^{N\times T}$ represent the input spike counts, where $\mathbf{x}_{i}^{t}$ denotes the spike count for neuron $i$ at time $t$. Let $\mathbf{b}\in\mathcal{B}=\mathbb{R}^{B\times T}$ be the corresponding behavior signal, with $\mathbf{b}_{i}^{t}$ denoting the behavioral signal at time $t$. Without loss of generality, we assume that behavioral signals are temporally continuous and have the same number of time steps as the neural activity. [^1] BLEND can be readily extended to accommodate scenarios wherein behavioral information is discrete or non-temporal in nature. As illustrated in Fig. 1, the model $f_{\theta}$ aims to reconstruct the randomly masked neural activity solely based on the unmasked portions of $\mathbf{x}$ or together with the auxiliary behavior signal $\mathbf{b}$. The optimization objective for masked time-series modeling (MTM) can be formulated as follows:
@@ 3.2.1 Privileged knowledge distillation with behavior as guidance
As is illustrated in Fig. 2(a), BLEND employs a teacher-student architecture to realize our goal: by leveraging the privileged information $\mathbf{b}$ at the training phase, to learn an NVM model for inference phase that outperforms those built on the regular feature $\mathbf{x}$ alone. This procedure is divided into two sequential stages:
@@
where $\alpha\in(0,1)$ denotes the mixing ratio between the MTM loss and the distillation loss. The entire optimization procedure of the proposed BLEND framework is presented in Alg. 1. After training the student model with privileged knowledge distillation for MTM, it is applied to downstream tasks where only neural activity recordings are available for inference (illustrated in Fig. 2(b)).
@@ 4 Experiments
As shown in Fig. 2(b), this work includes two benchmarks for evaluating our proposed BLEND framework. The first is a public benchmark for neural latent dynamics model evaluation from Pei et al. (2021), named Neural Latents Benchmark’21 (NLB’21). The second is a recent, public multi-modal neural dataset from Bugeon et al. (2022), which contains calcium imaging recordings of neural population activity as well as single-cell spatial transcriptomics of the recorded tissue. Details of these two benchmarks, tasks, and included baselines are introduced in Section 4.1 and 4.2.
@@ 6 Conclusion and Discussion
Although the initial results of BLEND are promising, a few limitations remain to be addressed. In this work, we employ a straightforward approach to implement the behavior guidance by concatenating the neural activity and behavior information along the feature dimension (assuming they have the same length of time steps). This methodological choice is motivated by the relatively low-dimensional nature of behavioral signals, which precludes their use as an independent guidance source (e.g., as a query in the cross-attention mechanism). However, this simplistic guidance strategy may not fully capture the intricate relationship between neural activity and behavior. Future research should investigate advanced integration methods to better utilize the complex interactions between these data types, potentially improving the model’s ability to represent subtle neural patterns. Moreover, while the current study focuses on temporal behavioral signals that correspond directly with neural activity at each time step, future research should aim to extend BLEND to more generalized settings. This expansion could encompass non-temporal or discrete behavioral signals, thereby broadening the framework’s applicability to diverse neuroscientific domains. Such extensions could prove valuable in investigating sleep stages, categorizing social behaviors, and exploring other scenarios where behavior is not continuously paired with neural activity. This generalization would significantly enhance the versatility and utility of BLEND across a wider spectrum of neuroscience research paradigms.
@@ A.6.3 Model-specific Observations
For each distillation strategy employ in this study, the following key observations can be discerned from Figure 5:

### v2

@@ 1 Introduction
To this end, in this paper, we propose BLEND, the Behavior-guided neuraL population dynamics modElling framework via privileged kNowledge Distillation. As shown in Fig. 1, BLEND constitutes a student and a teacher model, in which the teacher trains on both behavior observations and neural activity recordings, then distills knowledge to guide the student which takes only neural activity as input. This ensures the student model can make predictions using only recorded neural activity during the deployment/inference stage, but also benefits from the guidance from behavior information during the training stage, making it more versatile for settings lacking behavioral data. Our main contributions are summarized as follows:
@@ 2 Related work
Neural dynamics modeling (NDM). NDM refers to a category of methods that aim to capture the dynamics of neural activity by using the activity recordings as inputs. A common and effective choice is the latent variable model (LVM), which leverages low-dimensional latent factors to interpret these dynamics. Various LVMs are developed, ranging from simple non-temporal models such as principal components analysis (PCA) (Cunningham & Yu, 2014) and its variants (Kobak et al., 2016), to linear dynamical systems (Macke et al., 2011; Gao et al., 2016), and to complex state space models like LFADS (Pandarinath et al., 2018). Since the advent of Transformer, its ability to capture temporal dependencies and long-range interactions makes it appropriate for neural data, with representative works such as NeuralDataTransformer (NDT) (Ye & Pandarinath, 2021), STNDT (Le & Shlizerman, 2022), and EIT (Liu et al., 2022). As shown in Fig. 1, methods in this category purely depend on neural activity recordings for neural population dynamics modeling yet ignore using the paired behavior information as guidance.
Neural dynamics modeling with behavior as a prior. In recent years, an emerging research direction has focused on jointly modeling neural population dynamics and behavioral signals. As illustrated in Fig. 1, these methods can be categorized into two types. The first inputs both behavior and neural activity and utilizes behavior signals to guide the learning of neural dynamics: pi-VAE (Zhou & Wei, 2020) considers the behavior variables as constraints for the construction of latent space of LVMs; CEBRA (Schneider et al., 2023b) utilizes behavior signals to construct contrastive learning samples for label-informed neural activity analysis. However, these approaches normally need to develop delicately designed modules or complex training strategies to achieve their goal. The second line of work aims to decompose neural activity into behavior-relevant and behavior-irrelevant dynamics and reconstruct both behavior signals and neural activity signals. To that end, a linear state-space model, PSID (Sani et al., 2021), is developed to decode population dynamics from motor brain regions. Nonlinear state-space models such as TNDM (Hurwitz et al., 2021) and SABLE (Jude et al., 2022) are proposed to further capture the nonlinear dynamics of neural activity and behavior observation. Nonetheless, these models assume a clear distinction between behaviorally relevant and irrelevant dynamics in input neural activity, which might not always be practical, potentially leading to oversimplification.
In contrast, our approach offers a model-agnostic learning paradigm that can be directly applied to existing LVM models without relying on strong assumptions, thereby circumventing the issues present in current LVM models that integrate behavior information as a prior.
@@
Figure 2: Illustration of the proposed BLEND framework, exemplified using neural spiking activity data. Left: Behavior-guided neural representation learning via privileged knowledge distillation. The teacher model is trained on a composite of neural activity and behavioral signals, subsequently distilling its knowledge to a student model that utilizes solely neural activity as input. Right: During the inference phase, the distilled student model is employed for neural population activity analysis and transcriptomic identity prediction tasks.
@@ 3 Methods
This section introduces the details of our proposed BLEND framework, i.e., the behavior-guided neural population dynamics modeling via privileged knowledge distillation. We start with problem formulation of behavior-guided neural dynamics modeling in Section 3.1, followed by an exposition of the BLEND algorithm in Section 3.2. Detailed exploration of the effectiveness of the BLEND framework can be found in Appendix A.8 and A.9, including empirical studies of implications of behavior guidance and how to choose distillation strategies for different models and data.
@@ 3.1 Problem Formulation
We concretize the learning problem in this study with neural spiking data. For each trial of the input neural activity, let $\mathbf{x}\in\mathcal{X}=\mathbb{N}^{N\times T}$ represent the input spike counts, where $\mathbf{x}_{i}^{t}$ denotes the spike count for neuron $i$ at time $t$. Let $\mathbf{b}\in\mathcal{B}=\mathbb{R}^{B\times T}$ be the corresponding behavior signal, with $\mathbf{b}_{i}^{t}$ denoting the behavioral signal at time $t$. Without loss of generality, we assume that behavioral signals are temporally continuous and have the same number of time steps as the neural activity. [^1] BLEND can be readily extended to accommodate scenarios wherein behavioral information is discrete or non-temporal in nature, see preliminary explorations in Appendix A.10. As illustrated in Fig. 1, the model $f_{\theta}$ aims to reconstruct the randomly masked neural activity solely based on the unmasked portions of $\mathbf{x}$ or together with the auxiliary behavior signal $\mathbf{b}$. The optimization objective for masked time-series modeling (MTM) can be formulated as follows:
@@ 3.2.1 Privileged knowledge distillation with behavior as guidance
As is illustrated in Fig. 2, BLEND employs a teacher-student architecture to realize our goal: by leveraging the privileged information $\mathbf{b}$ at the training phase, to learn an LVM model for inference phase that outperforms those built on the regular feature $\mathbf{x}$ alone. This procedure is divided into two sequential stages:
@@
where $\alpha\in(0,1)$ denotes the mixing ratio between the MTM loss and the distillation loss. The entire optimization procedure of the proposed BLEND framework is presented in Alg. 1. After training the student model with privileged knowledge distillation for MTM, it is applied to downstream tasks where only neural activity recordings are available for inference (shown in Fig. 2, right).
@@ 4 Experiments
As shown in Fig. 2, this work includes two benchmarks for evaluating our proposed BLEND framework. The first is a public benchmark for neural latent dynamics model evaluation from Pei et al. (2021), named Neural Latents Benchmark’21 (NLB’21). The second is a recent, public multi-modal neural dataset from Bugeon et al. (2022), which contains calcium imaging recordings of neural population activity as well as single-cell spatial transcriptomics of the recorded tissue. Details of these two benchmarks, tasks, and included baselines are introduced in Section 4.1 and 4.2.
@@ 6 Conclusion and Discussion
Although the initial results of BLEND are promising, a few limitations remain to be addressed. In this work, we employ a straightforward approach to implement the behavior guidance by concatenating the neural activity and behavior information along the feature dimension (assuming they have the same length of time steps). This methodological choice is motivated by the relatively low-dimensional nature of behavioral signals, which precludes their use as an independent guidance source (e.g., as a query in the cross-attention mechanism). However, this simplistic guidance strategy may not fully capture the intricate relationship between neural activity and behavior. Future research should investigate advanced integration methods to better utilize the complex interactions between these data types, potentially improving the model’s ability to represent subtle neural patterns. Moreover, while the current study focuses on temporal behavioral signals that correspond directly with neural activity at each time step, future research should aim to extend BLEND to more generalized settings, as we initially explored in Appendix A.10. This expansion could encompass non-temporal or discrete behavioral signals, thereby broadening the framework’s applicability to diverse neuroscientific domains. Such extensions could prove valuable in investigating sleep stages, categorizing social behaviors, and exploring other scenarios where behavior is not continuously paired with neural activity. This generalization would significantly enhance the versatility and utility of BLEND across a wider spectrum of neuroscience research paradigms.
@@ A.6.3 Model-specific Observations
For each distillation strategy employed in this study, the following key observations can be discerned from Figure 5:
@@
A.7 Cross-correlation Analysis
This subsection provides cross-correlation and mutual information analyses on the MC-Maze dataset to quantify the relationship between neural activity and behavioral variables.
The cross-correlation heatmap was generated by computing correlations between each neuron’s activity and behavioral variables (X and Y velocities) across different time lags. For each neuron, the data was flattened across trials and timepoints, detrended, and normalized before computing correlations using a Fast Fourier Transform (FFT) method. The resulting correlations were arranged into a matrix where each row represents a neuron, each column represents a time lag, and the color intensity indicates the correlation strength. As shown in Fig. 6, our analysis revealed robust temporal relationships between neural activity and movement kinematics. Nearly half of the recorded neurons ($47.4\%$) exhibited leading relationships with horizontal (X) velocity, while a slightly larger proportion ($54.7\%$) led vertical (Y) velocity components. Importantly, we observe median leads of 20ms (X) and 10ms (Y), indicating that neural activity and behavior exhibit complex, complementary dynamics.
The strength of these neural-behavioral relationships was confirmed by highly significant correlations for both movement components ($p<0.001$ for both X and Y velocities). This robust statistical coupling explains why behavior serves as an effective privileged information source in our BLEND knowledge distillation framework, i.e., it provides a reliable signal about the neural computations being performed. Crucially, we find that population-level mutual information (MI) is substantially higher than single-neuron MI, indicating that behavior captures coordinated neural dynamics that emerge at the population level. This difference demonstrates that behavior provides a window into population-level neural representations, which is a key motivation for our knowledge distillation approach.
Taken together, these analyses show that neural activity and behavior in the MC-Maze dataset are strongly coupled, with behavior offering complementary, contextual information about the underlying neural computations. By effectively leveraging these neural-behavioral relationships, our BLEND framework is able to achieve significant performance improvements in learning better neural population representations. The robust temporal structure and statistical significance of the observed cross-correlations further validate the importance of behavior as a privileged signal in our knowledge distillation approach.
Figure 6: Visualization of neural activity and behavior cross-correlation heatmap on MC-Maze dataset. Upper: cross-correlation between independent neurons and horizontal movement; Below: cross-correlation between independent neurons and vertical movement.
A.8 Empirical Study on Behavior Guidance
To understand why incorporating behavior signals leads to better neural activity reconstruction, we conducted detailed analyses of the relationships between behavioral and neural data (exemplified with MC-Maze dataset).
A.8.1 Neural Space Organization
We first examined how behavior signals influence the organization of neural representations using t-SNE visualization (Figure 7). The baseline model’s representations (shown in red) appear as scattered, disconnected clusters across the neural space. In contrast, BLEND’s representations (shown in blue) form more cohesive structures, suggesting that behavior signals help constrain neural activity patterns into meaningful manifolds. This improved organization likely contributes to BLEND’s superior reconstruction capability.
Figure 7: t-SNE visualization of neural representations showing the organization of neural space. Red points represent baseline model representations, while blue points show BLEND model representations. The more cohesive clustering of BLEND’s representations suggests better capture of underlying neural structure.
Figure 8: Behavior-neural analysis. (a) Neural activity patterns across different behavioral states show systematic variation. Different colors represent different behavior quartiles. (b) Box plots comparing reconstruction errors for neurons with low versus high behavior coupling, demonstrating improved reconstruction for behavior-coupled neurons.
A.8.2 Behavior-Dependent Neural Activity Patterns
We analyzed how neural activity patterns vary systematically with behavioral states (Figure 8, left). The behavioral states were determined through the behavior data, which consists of two continuous variables (x and y coordinates of hand position) over time. We used the first behavior variable (x coordinate) and divided it into quartiles using numpy’s percentile function: behavior_bins = percentile(behavior_data, [0, 25, 50, 75, 100]). This created four behavior states (1/4 through 4/4), representing different ranges of hand positions. For each state, we computed the mean neural activity pattern across all time points falling within that state. The resulting patterns show clear modulation of neural activity by behavioral state, with distinct activity profiles for different hand positions. Some neurons (e.g., indices 25, 50, and 125) show particularly strong behavior-dependent modulation.
A.8.3 Reconstruction Quality Analysis
We quantitatively assessed how behavior coupling influences reconstruction quality through a comprehensive analysis framework. First, we quantified behavior-neural coupling using mutual information (MI): $MI(N_{i},B)=\sum_{n,b}p(n,b)\log\frac{p(n,b)}{p(n)p(b)}$, where $N_{i}$ is the activity of neuron $i$ and $B$ is the behavior variable. We computed this for each neuron using scikit-learn’s mutual_info_regression function. Neurons were then classified into high coupling (MI $>$ median) and low coupling (MI $\leq$ median) groups.
For each neuron $i$, we computed the reconstruction error as $E_{i}=\frac{1}{T}\sum_{t=1}^{T}(\hat{r}_{i}^{t}-r_{i}^{t})^{2}$, where $\hat{r}_{i}^{t}$ is the predicted firing rate and $r_{i}^{t}$ is the true firing rate at time $t$. The box plots (Figure 8, right) show the distribution of these errors for both groups. Statistical analysis revealed significant differences between high and low coupling groups (Wilcoxon rank-sum test: $p<0.001$), with high-coupling neurons showing consistently lower reconstruction errors.
These analyses collectively demonstrate that BLEND’s improved performance stems from its ability to leverage inherent structure in behavior-neural relationships. The systematic variation of neural activity with behavior and the improved reconstruction for behavior-coupled neurons provide strong empirical evidence for the effectiveness of incorporating behavior signals in neural activity reconstruction.
A.9 Distillation Strategy Evaluation on Synthetic Data
This subsection provides a detailed exploration and a comprehensive understanding of when and why different strategies excel. We conduct additional experiments using carefully designed synthetic datasets that mirror the structure of real neural recordings while allowing precise control over neural-behavioral relationships.
We created three types of synthetic datasets with distinct characteristics.
  • Simple: The first type implements simple linear relationships between neural activity and behavior, designed to test basic knowledge transfer. In this dataset, neural activity is directly derived from sinusoidal behavioral trajectories through a linear transformation. This creates a clear and interpretable relationship between behavior and neural responses, with added Poisson noise to simulate realistic spike counts.
  • Hierarchical: The second type features hierarchical relationships where different neuron groups encode behaviors at varying timescales, simulating the layered processing often observed in neural circuits. The hierarchical dataset introduces temporal complexity by dividing neurons into three functional groups, each processing behavior at different timescales. Fast neurons respond to immediate behavior, medium neurons integrate over 5 timepoints, and slow neurons average over 10 timepoints, creating a rich temporal hierarchy that mirrors the multi-timescale processing observed in real neural systems.
  • Complex: The third type incorporates complex population-level correlations with temporal dependencies, mimicking the intricate dynamics found in real neural populations. The complex population dataset represents the most sophisticated model, organizing neurons into five correlated assemblies with explicit population-level structure. This dataset features decaying sinusoidal behavior patterns and maintains specific correlation structures (0.3 within groups) through multivariate normal noise, creating realistic population dynamics.
Methods Simple Hierarchical Complex
LFADS-Hard-Distill $0.985$ $0.976$ $0.182$
NDT-Hard-Distill $0.934$ $0.926$ $0.147$
NDT-Soft-Distill $0.927$ $0.932$ $0.146$
NDT-Feature-Distill $0.926$ $0.940$ $0.139$
NDT-Correlation-Distill $0.935$ $0.944$ $0.180$
Table 4: Results showing variants of BLEND framework on three synthetic datasets with distinct characteristics. Vel-$R^{2}$ is used as the measurement metric.
All three datasets share common dimensions (2,869 trials, 140 timepoints, 182 neurons, 2 behavioral variables) and are systematically divided into train/eval splits and held-in/held-out neurons, enabling rigorous testing of neural analysis methods.
Our experiments with these synthetic datasets revealed clear patterns that explain the performance differences observed in Table 4. LFADS consistently performed better with Hard Distillation across all synthetic cases, particularly excelling with simple linear relationships (around 6% improvement over NDT-based variants in Vel-$R^{2}$). This aligns with our understanding that LFADS’s RNN-based architecture, with its inherent temporal continuity and built-in constraints, benefits most from direct knowledge transfer. The architecture’s strong internal regularization makes it well-suited for precise, deterministic knowledge transfer through hard distillation.
Conversely, NDT showed superior performance with Correlation Distillation, especially in cases with complex population-level correlations (around 20% improvement over other NDT-based variants). This finding stems from NDT’s transformer-based architecture, which processes information in parallel and lacks inherent temporal constraints. The correlation distillation strategy provides valuable structural guidance that complements NDT’s flexible architecture, helping it maintain important population-level relationships in the learned representations.
Intuitions and guidelines: These insights from synthetic data experiments provide concrete guidelines for strategy selection in practice. For architectures with strong internal regularization like LFADS, Hard Distillation offers the most direct and effective knowledge transfer. For more flexible architectures like NDT, Correlation Distillation helps maintain complex population dynamics and temporal relationships. Soft Distillation, serving as a robust middle-ground option, can be beneficial for datasets where subtle variations in neural responses need to be preserved, though it may not fully capture complex population-level dynamics. Feature Distillation, while generally showing more modest improvements, is particularly effective when the neural data exhibits a clear hierarchical structure, as it can capture relationships across different levels of neural representation. This understanding not only explains the performance patterns in our original results but also offers a principled approach to selecting distillation strategies for future applications of BLEND to different neural datasets and architectures.
A.10 Extending BLEND to Discrete or Non-temporal Behavior Labels
To verify the capability of BLEND with discrete or non-temporal behavior observations, we extend BLEND on the DMFC-RGS dataset from NLB’21 Benchmark Pei et al. (2021), which contains recordings from dorso-medial frontal cortex (DMFC) during a cognitive timing task known as Ready-Set-Go (RSG). The behavioral variables in this task are distinctly discrete in nature. The task incorporates five key behavioral features, represented as binary or categorical variables: ’is_eye’ indicating whether the response was an eye movement (1) or joystick movement (0), ’theta’ specifying the categorical direction of movement (left or right), and ’is_short’ denoting whether the trial used the Short prior (1) or Long prior (0). The remaining variables ’ts’ and ’tp’ represent the sample and produced time intervals respectively. These discrete behavioral choices create a structured experimental design with clearly defined conditions.
Given that this dataset has discrete and non-temporal behavior labels, we align the neural activity and behavior by introducing a learnable temporal position encoding for discrete behavior variables across the length of neural recording. We compare NDT-Hard-Distill with the original NDT model on this dataset, and this simple extension shows improvement in both Co-Bps ($0.127$ achieved by NDT-Hard-Distill and $0.118$ achieved by NDT) and PSTH-$R^{2}$ ($0.444$ achieved by NDT-Hard-Distill and $0.338$ achieved by NDT).
These promising results demonstrate the effectiveness of BLEND on both continuous/discrete and temporal/non-temporal behavior observations, showcasing a flexible solution we provide for learning better neural population dynamics.